Finding the equation of a line given the following points:
(-2,-5) (-8,-6)

Answer by rfer(16322)

m=(-6--5)/(-8--2)
m=-1/-6
m=1/6
---------
-6=1/6(-8)+b
-6=-8/6+b
-6+8/6=b
-36/6+8/6=b
-28/6=b
-4.7=b
----------------
y=1/6x-4.7